OpenStudy (anonymous):

6^2-(1/2)^3/7/8+5/12

OpenStudy (anonymous):

\[6^2-\frac{(\frac{1}{2})^3}{\frac{7}{8}}+\frac{5}{12}\]

OpenStudy (anonymous):

\[36-(\frac{1}{2})^3*\frac{8}{7}+\frac{5}{12}\]

OpenStudy (anonymous):

\[36-\frac{1}{8}*\frac{8}{7}+\frac{5}{12}\]

OpenStudy (anonymous):

\[36-\frac{1}{7}+\frac{5}{12}\]

OpenStudy (anonymous):

\[\frac{36}{1}*\frac{12*7}{12*7}-\frac{1}{7}*\frac{12}{12}+\frac{5}{12}*\frac{7}{7}\]

OpenStudy (anonymous):

\[\frac{3024-12+35}{84}\]

OpenStudy (anonymous):

\[\frac{3047}{84}\]

OpenStudy (anonymous):

does it matter that the equation is 6^2-1/2^3 is divided by 7/8+5/12? i was unable topost it as it is wriiten in the problem. it s not 6^2 and then minus the fraction 1/2^3/7/8 and then plus 5/12. how do you post a fraction like you did? 3047/84?

OpenStudy (anonymous):

you see equation down below go there and type frac{something}{something}

OpenStudy (anonymous):

you can also a fraction in a fraction if you want

OpenStudy (anonymous):

\[\left(\begin{matrix}6^2-1/2^3 \\ 18+5/12 ?\end{matrix}\right)\]

OpenStudy (anonymous):

\[\frac{6^2-(\frac{1}{2})^3}{18+\frac{5}{12}}\]?

OpenStudy (anonymous):

i ouldnt get the line in middle..lol but thats not right either..its posed to be 7/8..not 18 \[\left[\begin{matrix}6^2-1/2^3 \\ 7/8+5/12\end{matrix}\right]\]

OpenStudy (anonymous):

\[\frac{36-\frac{1}{8}}{\frac{7}{8}+\frac{5}{12}}\]

OpenStudy (anonymous):

\[\frac{\frac{36}{1}*\frac{8}{8}-\frac{1}{8}}{\frac{7}{8}*\frac{12}{12}+\frac{5}{12}*\frac{8}{8}}\]

OpenStudy (anonymous):

\[\frac{\frac{287}{8}}{\frac{84+40}{96}}=\frac{287}{8}*\frac{96}{114}\]

OpenStudy (anonymous):

i will leave the simplification to you im going to go eat

OpenStudy (anonymous):

ok thanks i will figure it out